OpenStudy (jellybot23):

I know this is more chemistry, but I am confused on what to do mathematically! HELP What is the pH of a 0.085 M solution of nitrous acid (HNO2) that has a Ka of 4.5 × 10-4?

OpenStudy (jellybot23):

@Michele_Laino

Vocaloid (vocaloid):

first step: is nitrous a strong acid or weak acid?

OpenStudy (jellybot23):

a weak acid, correct?

Vocaloid (vocaloid):

good we can now write out our acid dissociation equation

Vocaloid (vocaloid):

Ka = [H+][NO2-]/[HNO2] we let x = the concentration of H+ giving us let C = the original concentration Ka = x * x / (C-x) now we plug in our Ka and C values to solve for x

Vocaloid (vocaloid):

4.5*10^(-4) = x^2/(0.085 - x)

Vocaloid (vocaloid):

shortcut method: we can assume that the dissociated x is very small, so 0.085 - x is almost 0.085

Vocaloid (vocaloid):

4.5*10^(-4) = x^2/0.085 solve for x

OpenStudy (jellybot23):

I tried that before, but got a ridiculously small number. I get 3.825X10^-5 = x^2 and then the square root of that is around .006184

OpenStudy (michele_laino):

correct!

Vocaloid (vocaloid):

it's supposed to be small next step is to find -log(x)

Vocaloid (vocaloid):

where x = .006184

Vocaloid (vocaloid):

(remember, x is not the pH. x is the H+ concentration)

OpenStudy (jellybot23):

Ohhhh That's where I was getting confused. so when I do -log(.006184) I get about 2.21!

Vocaloid (vocaloid):

good! that's correct

OpenStudy (jellybot23):

Yay! Thank you so much!